OpenStudy (anonymous):

How would I solve the improper integral 1/(sqrt(x^2+1) from 0 to infinity

OpenStudy (anonymous):

do you know the anti derivative of \(\frac{1}{\sqrt{x^2+1}}\)?

OpenStudy (anonymous):

yes @satellite73 it's going to be arctan

OpenStudy (anonymous):

what the hell is this???

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

this business you have copied and pasted above that has nothing to do with this problem?

OpenStudy (anonymous):

i miss my old delete button, need to get it back.

OpenStudy (ash2326):

@Rohangrr last question also you copied and pasted from somewhere else!!! Please use your own work here

OpenStudy (anonymous):

@ash2326 i dont copy dam!! make that sure

OpenStudy (anonymous):

I just know that improper integration of 1/(sqrt(x^2+1) is \[\ln(x+\sqrt{x^2 + 1}) \] Now apply upper limit - lower limit and get your answer

OpenStudy (anonymous):

u dont know the ans... thats improper integral if i dont get to know i have ms word i solve in my copy n write there!! @ash2326

OpenStudy (ash2326):

Ok then will you explain it @Rohangrr

OpenStudy (anonymous):

@Rohangrr you are making it difficult to answer this question. the material you copied and pasted has nothing to do with the question asked

OpenStudy (anonymous):

sorry i had written some thing else!! @satellite73 yep sorry dude

OpenStudy (anonymous):

neither did your last 3 answers so maybe it is time for you to take a break

OpenStudy (anonymous):

@shivam_bhalla that's what I got for the improper integral now all I have to do is evaluate at the limits thank you :)

OpenStudy (anonymous):

lets repost and we can do this problem without much difficutly

OpenStudy (anonymous):

:) Welcome

OpenStudy (anonymous):

I will post my next question shortly I am doing exam prep and I just needed help with a couple series and improper integral questions thanks for your assistance

OpenStudy (ash2326):

@satellite73 We can evaluate the indefinite integral but the limits here 0 to \(\infty\). I think the integral doesn't converge for these limits

OpenStudy (anonymous):

@Rohangrr Wolfram much?

OpenStudy (anonymous):

no it does not converge because as x gets large, so does \(\ln(x)\)

OpenStudy (anonymous):

@Romero dont tungsten much!

OpenStudy (anonymous):

@Rohangrr Did you just wolframed wolfram ?